# LLM-SQL-Solver: Can LLMs Determine SQL Equivalence?

### 16 Dec 2023 v1

@@ 1 Introduction
Structured Query Language (SQL) it is the standard programming language to query data systems and it is one of the most popular programming language  [^1] https://www.stackscale.com/blog/most-popular-programming-languages/. Since SQL is not a procedural language, one may write different SQL queries that are semantically equivalent. As a result, determining the equivalence of SQL queries is a fundamental task in many practical important applications ranging from query optimizers in database systems Pirahesh et al. (1992); Bai et al. (2023); Zhou et al. (2021) to evaluating the translation from user text-to-SQL Zelle and Mooney (1996); Zhong et al. (2020b); Khatry et al. (2023). However, it has been shown by the data management research community that checking the equivalence of two SQL queries is in general undecidable Abiteboul et al. (1995). While there are theoretical works to check equivalence under different assumptions Chandra and Merlin (1977); Sagiv and Yannakakis (1980); Cohen et al. (1999); Chu et al. (2018) and implemented prototypes such as Cosette Chu et al. (2017), they only support limited SQL operations and such projects are no longer maintained due to lack of experts. For instance, Cosette does not support commonly found operations like EXISTS, IN, queries with constant expressions, etc.; and Cosette is no longer maintained due to the lack of staff. Query optimizers and query processing engines in data systems still primarily rely on heuristic query rewriting rules Begoli et al. (2018). Exploring and writing these rules correctly require strong domain expertise and need many engineering hours.
Moreover, due to the difficulty in evaluating the equivalence of two SQL queries, Text-to-SQL literature Zhong et al. (2017); Yu et al. (2018); Lee et al. (2021); Li et al. (2023) primarily use exact match (exact SQL string match) or execution match (compare the execution outputs) as the evaluation metrics Zhong et al. (2020a). However, exact string match may overlook many equivalent SQL queries (there are many ways different ways to write the same SQL query) and execution match may misclassify different SQL queries as equivalent due to no counterexamples are contained in the underlining database instance.
Recent advancements in large language models (LLMs) have shown great capability in following instructions and in generating decisions that align with human judgement Bubeck et al. (2023); Chiang and Lee (2023); Yan et al. (2023a). For instance, It has been shown that GPT models are capable of performing first order logic reasoning  Saparov and He (2022), detect implicit contents Huang et al. (2023); Zhang et al. (2023b), and refining the invisible logic Feng et al. (2023). In this paper, we investigate whether LLM can act as a judge to determine the equivalence of SQL queries. Hence, in this paper, we would like to mainly answer the following questions:
@@ 4.3.2 Human Evaluation on Samples
Equivalence Evaluated Manually To evaluate the equivalence between two SQL queries, we use the human preference approach. We manually check the equivalence of the sampled questions where LLMs and execution match disagree. In human evaluation, we have also differentiated between the semantic equivalence and relaxed equivalence (in essence, semantic equivalence is objective and relaxed equivalence is subjective),
Quality Control To ensure the quality of human judgement, we have three independent SQL experts judging the equivalence and we use the majority vote to determine the label for each pair of SQL queries.
@@
DIN 7/10 $\uparrow$
DAIL 5/10 -
GPT 3.5 turbo Relaxed Chat 6/10 $\uparrow$
@@
DAIL 6/10 $\uparrow$
[…]
DIN 7/10 $\uparrow$
DAIL 9/10 $\uparrow$
@@
In Table 4, GPT 3.5 turbo aligned with human experts on 2 out of 10 questions and 4 out of 10 questions for Chat and DIN respectively. Diving deep into the questions which GPT 3.5 turbo wrongly classified the SQL queries, we find that there are common reasons for the wrong classifications: Distinct entries vs duplicated entries. GPT 3.5 turbo often failed to consider the difference between selecting unique rows and selecting rows that may be duplicated; Subtle difference in column names. When the gold SQLs and generated SQLs share very similar SQL structure but contains subtle differences in the columns used such as group by id v.s. group by student_id, GPT 3.5 turbo may fail to notice the small difference; Hallucination. GPT 3.5 turbo may returns wrong results from executing the SQL queries on its own created simple database instance and sometimes when the outputs are indeed identical, it may still state that the outputs are different. This hallucination snowballing effect Zhang et al. (2023a) is one of the top reason why GPT 3.5 turbo perform worse compared to GPT 4 turbo.
When GPT 4 turbo is used as the judge, GPT 4 turbo constantly produced more accurate or comparable results to execution match, with at least 5 out of 10 questions for DAIL and at most 10 out of 10 questions for Chat. One interesting observation is that GPT 4 turbo correctly identified the difference between order by asc/desc limit 1 v.s. where col = select min/max(col). The former will always return one row, as limit 1 is used. The latter, however, may return many rows when these rows share the same min or max value. Execution match often failed to capture this difference. Although GPT 4 turbo judgment is much better than GPT 3.5 turbo, GPT 4 turbo may still miss on the small differences such as similar but different column names (song_name v.s. name), overlook the fact that extra columns are selected, and hallucinate. As shown in Figure 2, we looked into how well does LLMs executed and reasoned about the SQL queries on its own generated simple database. We find that GPT 4 turbo has much higher correct operation rates and hence much better response.
@@
Relaxed Equivalence When using the explain and compare prompt for determining the relaxed equivalence, both GPT 3.5 turbo and GPT 4 turbo consistently better aligned with human judgment compared with execution match, as shown in Table 4. GPT 3.5 turbo aligned with human preference on at least 6 out of 10 questions and GPT 4 turbo aligned with human preference on at least 7 out of 10 questions. Recall, the relaxed notion mimics the opinion of the users of NL2SQL services. The positive results on the alignment between LLMs judgments with human preferences indicate optimizing for LLMs judgments accuracy is a better optimization target than execution match accuracy.
The main reasons for why LLMs judgements are better aligned with human preference than execution match is that LLMs can zoom into the logic and determine whether the two SQL queries use the same exact logic and contain the same truth. For instance, order by asc/desc limit 1 v.s. where col = select min/max(col) may return different or the same outputs depending on the underlying database instance. Since execution match strictly checks the outputs, it may sometimes return equivalent and other times return not equivalent. However, based on the relaxed equivalence definition, order by asc/desc limit 1 is the same as where col = select min/max(col) since they both are selecting the rows with minimal or maximal value.
@@ 5 Discussion and Limitation
There are some limitations or areas of improvement to this work. First, we only evaluated the capability of Large Language Models of OpenAI. There are more other models such as Llama2 7B Touvron et al. (2023), Vicuna Chiang et al. (2023), Alpaca Taori et al. (2023a) which also showcase strong understanding and reasoning for logical tasks Chen et al. (2021); Yue et al. (2023); Yan et al. (2023b). Nevertheless, we believe that GPT 3.5 turbo and GPT 4 turbo exemplify the current status of GPT model family’s capabilities. In addition, in the manually evaluating the equivalence under two different views (semantic and relaxed), it is done over the sampled questions. Since there are more than two thousand of SQL queries per data source and equivalence view, manually checking all of them are arguably infeasible. In fact, even checking a pair of SQL queries for semantic view takes around 5 minutes for experts in SQL. Although the questions are sampled, we observe a consistent comparisons and trends between LLMs’ judgements and execution matches (such as GPT 4 turbo’s responses are consistently better aligned with human preferences).
@@ 6 Conclusion
In this paper, we studies and answer two main questions: i) Can current LLMs be used to help data engineers in checking semantic equivalences between SQL queries and ii) should we use LLMs to judge the quality of generated SQL queries for text-to-SQL benchmarks. Through extensive evaluations and manually examining the quality of the answers based on the alignment with human preferences, we find current LLMs (especially GPT 4 turbo), using our proposed prompt techniques, can assists engineers in determining semantic equivalent queries, but may yield incorrect responses. In addition, when comparing the generated SQL query with the gold SQL query, LLMs’ responses achieve high alignments with human preferences. Moreover using LLMs to determine relaxed SQL equivalence avoid the construction of sophisticated databases and no data stored need to be accessed. To be best of our knowledge, this is the first work studying the ability of LLMs in determining SQL equivalence. As the learning ability in LLMs continuously to advance, we believe that using these models to check SQL equivalence is a promising direction.

### 17 Jan 2024 v2

@@ 1 Introduction
Structured Query Language (SQL) it is the standard programming language to query data systems and it is one of the most popular programming language  [^1] https://www.stackscale.com/blog/most-popular-programming-languages/. Since SQL is not a procedural language, one may write different SQL queries that are semantically equivalent. As a result, determining the equivalence of SQL queries is a fundamental task in many practical important applications ranging from query optimizers in database systems Pirahesh et al. (1992); Bai et al. (2023); Zhou et al. (2021); Samwel et al. (2018) to evaluating the translation from user text-to-SQL Zelle and Mooney (1996); Zhong et al. (2020b); Khatry et al. (2023); Floratou et al. (2024). However, it has been shown by the data management research community that checking the equivalence of two SQL queries is in general undecidable Abiteboul et al. (1995). While there are theoretical works to check equivalence under different assumptions Chandra and Merlin (1977); Sagiv and Yannakakis (1980); Cohen et al. (1999); Chu et al. (2018) and implemented prototypes such as Cosette Chu et al. (2017), they only support limited SQL operations and such projects are no longer maintained due to lack of experts. For instance, Cosette does not support commonly found operations like EXISTS, IN, queries with constant expressions, etc.; and Cosette is no longer maintained due to the lack of staff. Query optimizers and query processing engines in data systems still primarily rely on heuristic query rewriting rules Begoli et al. (2018). Exploring and writing these rules correctly require strong domain expertise and need many engineering hours.
Moreover, due to the difficulty in evaluating the equivalence of two SQL queries, Text-to-SQL literature Zhong et al. (2017); Yu et al. (2018); Wang et al. (2019); Lee et al. (2021); Li et al. (2023) primarily use exact match (exact SQL string match) or execution match (compare the execution outputs) as the evaluation metrics Zhong et al. (2020a). However, exact string match may overlook many equivalent SQL queries (there are many ways different ways to write the same SQL query) and execution match may misclassify different SQL queries as equivalent due to no counterexamples are contained in the underlining database instance.
Recent advancements in large language models (LLMs) have shown great capability in following instructions and in generating decisions that align with human judgement Bubeck et al. (2023); Chiang and Lee (2023); Yan et al. (2023a); Zhang et al. (2023b). For instance, It has been shown that GPT models are capable of performing first order logic reasoning  Saparov and He (2022), detect implicit contents Huang et al. (2023); Zhang et al. (2023c), and refining the invisible logic Feng et al. (2023). In this paper, we investigate whether LLM can act as a judge to determine the equivalence of SQL queries. Hence, in this paper, we would like to mainly answer the following questions:
@@ 4.3.2 Human Evaluation on Samples
Equivalence Evaluated Manually To evaluate the equivalence between two SQL queries, we use the human preference approach. Three of the authors manually check the equivalence of the sampled questions where LLMs and execution match disagree. In addition, we have also differentiated the equivalence between the semantic notion and relaxed notion (in essence, the former is objective and the latter is subjective). To ensure the quality of human judgement, the annotations are conducted independently, and we use the majority vote to determine the label for each pair of SQL queries (i.e., each pair of SQL queries are examined by three times independently).
@@
DAIL 5/10 -
[…]
DIN 6/10 $\uparrow$
DAIL 5/10 -
GPT 3.5 turbo Relaxed Chat 6/10 $\uparrow$
DIN 5/10 $\uparrow$
@@
DAIL 7/10 $\uparrow$
@@
In Table 4, GPT 3.5 turbo aligned with human experts in the range of two to five out of ten. Diving deep into the questions which GPT 3.5 turbo wrongly classified the SQL queries, we find that there are common reasons for the wrong classifications: Distinct entries vs duplicated entries. GPT 3.5 turbo often failed to consider the difference between selecting unique rows and selecting rows that may be duplicated; Subtle difference in column names. When the gold SQLs and generated SQLs share very similar SQL structure but contains subtle differences in the columns used such as group by id v.s. group by student_id, GPT 3.5 turbo may fail to notice the small difference; Hallucination. GPT 3.5 turbo may returns wrong results from executing the SQL queries on its own created simple database instance and sometimes when the outputs are indeed identical, it may still state that the outputs are different. This hallucination snowballing effect Zhang et al. (2023a) is one of the top reason why GPT 3.5 turbo perform worse compared to GPT 4 turbo.
When GPT 4 turbo is used as the judge, GPT 4 turbo constantly produced more accurate or comparable results to execution match, with at least 5 out of 10 questions for DAIL and at most 10 out of 10 questions for Chat. One interesting observation is that GPT 4 turbo correctly identified the difference between order by asc/desc limit 1 v.s. where col = select min/max(col). The former will always return one row, as limit 1 is used. The latter, however, may return many rows when these rows share the same min or max value. Execution match often failed to capture this difference. Although GPT 4 turbo judgment is much better than GPT 3.5 turbo, GPT 4 turbo may still miss on the small differences such as similar but different column names (such as song_name v.s. name), overlook the fact that extra columns are selected, and also sometimes hallucinate. As shown in Figure 2, we looked into how well does LLMs executed and reasoned about the SQL queries on its own generated simple database. We find that GPT 4 turbo has much higher correct operation rates which leads much better response comapred with the weaker model.
@@
Relaxed Equivalence When using the explain and compare prompt for determining the relaxed equivalence, both GPT 3.5 turbo and GPT 4 turbo are consistently better aligned with human judgment or at par compared to execution match, as shown in Table 4. Recall, the relaxed notion aims to illustrate the degree to which humans perceive two queries as being similar in their meaning, and mimics the opinion of the users of NL2SQL services. The positive results on the alignment between LLMs judgments with human preferences indicate optimizing for LLMs judgments accuracy is a better optimization target than execution match accuracy.
The main reasons for why LLMs judgements are better aligned with human preference than execution match is that LLMs can zoom into the query logic and determine whether the two SQL queries share the same logic and contain the same truth. Consider the following example, order by col asc/desc limit 1 v.s. where col = select min/max(col). They may return different or the same outputs after execution (depends on the underlying database instance). Since execution match strictly checks the outputs, it may sometimes return equivalent and other times return not equivalent. However, LLMs can correctly identify that these two queries are selecting the rows with minimal or maximal value and these two queries are equivalent under the relaxed notion,
@@ 5 Discussion and Limitation
There are some limitations or areas of improvement to this work. First, we only evaluated the capability of Large Language Models of OpenAI. There are more other models such as Llama2 7B Touvron et al. (2023), Vicuna Chiang et al. (2023), Alpaca Taori et al. (2023a) which also showcase strong understanding and reasoning for logical tasks Chen et al. (2021); Yue et al. (2023); Yan et al. (2023b). Nevertheless, we believe that GPT 3.5 turbo and GPT 4 turbo exemplify the current status of GPT model family’s capabilities. In addition, in the manually evaluating the equivalence under two different views (semantic and relaxed), it is done over the sampled questions. Since there are more than two thousand of SQL queries per data source and equivalence view, manually checking all of them are arguably infeasible. In fact, even checking a pair of SQL queries for semantic view takes around 10 minutes for experts in SQL. Although the questions are sampled, we observe a consistent trends between LLMs’ judgements and execution matches (such as GPT 4 turbo’s responses are consistently better aligned with human preferences).
@@ 6 Conclusion
In this paper, we studies and answer two main questions: i) Can current LLMs be used to help data engineers in checking semantic equivalences between SQL queries and ii) should we use LLMs to judge the quality of generated SQL queries for text-to-SQL benchmarks. Through extensive evaluations and manually examining the quality of the answers based on the alignment with human preferences, we find current LLMs (especially GPT 4 turbo), using our proposed prompt techniques, can assists engineers in determining semantic equivalent queries, but may yield incorrect responses. On the other hand, when comparing the generated SQL query with the gold SQL query with relaxed notion of equivalence, LLMs’ responses constantly achieve high alignments with human preferences. Moreover, another advantage of using LLMs to determine SQL equivalence is that it can avoid the construction of sophisticated databases, and no data stored need to be accessed. To the best of our knowledge, this is the first work studying the ability of LLMs in determining SQL equivalence. As the learning ability in LLMs continuously to advance, we believe that using these models to check SQL equivalence is a promising direction.